OpenStudy (anonymous):

A triangle has one side that measures x, and the other two sides each measure 4 inches less than x. The perimeter is 19 inches. What is the measure of x? (A) 2 (B) 3 (C) 5 (D) 9

OpenStudy (anonymous):

x + 2(x-4) = 19 3x = 19 + 8 = 27 x = 9